OpenStudy (camerondoherty):

A system of equations is shown below: 3x + 8y = 12 2x + 2y = 3 Create an equivalent system of equations by replacing one equation with the sum of that equation and a multiple of the other. Show the steps to do this. Show that the equivalent system has the same solution as the original system of equations.

OpenStudy (anonymous):

Do you understand the question? @camerondoherty

OpenStudy (dan815):

you can chose other multiples

OpenStudy (dan815):

3x+8y=12 2x+2y=3 ------------- 5x+10y=15 <-------- get by adding the 2 given equations 4x+4y=6 <---------- multiple by 2 of the 2

OpenStudy (dan815):

multiple by 2 of the 2nd given eqn

OpenStudy (dan815):

Those 2 new equations under the separation '------' can be your new new 2 eqns

OpenStudy (dan815):

5x+10y=15 4x+4y=6 ---------------- solve this system and see what x and y have to be, then you can plug it back into the first 2 given equations, and show that this x and y value satisfy those equations, so this means these 2 sets of system of equations are equivalent.

OpenStudy (camerondoherty):

Thanks @dan815 and sowwy i was eating lol